1 Dihedral groups

The dihedral group D_{n} is the symmetry group of a regular n-sided polygon. Generated by a rotation r of 2π/n radians and by a mirroring operation s, there are 2n elements in the group D_{n}:

D_{n} = {e, r, r^{2}, ..., r^{n−1}, s, rs, ..., r^{n−1}s}  (1)